# Magnetic fields limit the mass of Population III stars even before the onset of protostellar radiation feedback

### v1

@@ 1 Introduction
However, radiation-magnetohydrodynamics simulations in the era of Pop III star formation remain largely absent. Both magnetic fields and protostellar radiation feedback are critical ingredients that influence (massive) star formation across all metallicities (e.g., Tanaka et al., 2018; Chon et al., 2021, 2024; Sharda & Krumholz, 2022). Therefore, conclusions from prior numerical work aimed at deriving the masses of Pop III stars are likely subject to major uncertainties since they either exclude magnetic fields (e.g., Hosokawa et al., 2011, 2016; Sugimura et al., 2020; Jaura et al., 2022) or radiation feedback (e.g., Turk et al., 2012; Sharda et al., 2020; Prole et al., 2022; Saad et al., 2022). In this work, we use the POPSICLE project (Sharda et al. in prep.) to extend the recent RMHD simulations of Pop III star formation by Sharda & Menon (2024) to include far-UV (FUV) molecule-dissociating radiation feedback in addition to extreme-UV (EUV) ionizing feedback. Our aim in this Letter is to show that magnetic fields significantly limit the mass growth of massive Pop III stars, even before radiative feedback becomes dominant. We arrange the remainder of the paper as follows: Section 2 summarizes the setup we use to develop and run the simulations, and Section 3 describes the results. In Section 4, we look at the main sequence evolution of simulated stars using Modules for Experiments in Stellar Astrophysics (MESA, Paxton et al., 2011, 2013, 2015, 2018, 2019). Finally, we summarize in Section 5.
@@ 2 POPSICLE Simulations
We briefly describe the POPSICLE [^1] POPulation II/III Simulations Including Chemistry, Luminosity, and Electromagnetism. project setup employed in this work, and point the reader to Sharda & Menon (2024) for details of the numerical implementation. We use a custom version of the adaptive mesh refinement (AMR) code FLASH (Fryxell et al., 2000; Dubey et al., 2008) which employs an approximate Riemann solver for magnetohydrodynamics (Waagan, 2009; Waagan et al., 2011). We include non-equilibrium primordial chemistry (with deuterium) from the KROME astrochemistry package (Grassi et al., 2014). We use the VETTAM radiation hydrodynamics scheme, which uses a non-local variable Eddington tensor (VET) closure obtained with a ray-trace solve to close the radiation moment equations (Menon et al., 2022). We use the GENEVA Pop III protostellar model grid to evolve radiative properties of the protostars as a function of their mass and accretion rates (Haemmerlé et al., 2016, 2018). The protostars are represented by sink particles in the simulation, following the criteria described in Federrath et al. (2010, 2011).
Population III protostars contract and produce significant radiation once accretion rates drop below $0.01\,\rm{M_{\odot}\,yr^{-1}}$ (Omukai & Palla, 2001, 2003). We consider the ionization of H and H2 due to extreme-UV (EUV) photons released from the protostar(s) with energies upwards of $13.6\,\rm{eV}$ (upwards of $15.2\,\rm{eV}$ for H2) released from the stars that form in the simulations. The key physics we add to the simulations by Sharda & Menon (2024) is the dissociation of H2 by FUV photons in the Lyman-Werner (LW) band, between $11.2-13.6\,\rm{eV}$, which can significantly affect the thermodynamic state of dense gas in the vicinity of the stars. Importantly, we consider both self-shielding of H2 as well as cross shielding by H (Wolcott-Green & Haiman, 2011); the latter has been ignored in previous works. We adopt the fitting functions for self- and cross-shielding of H2 from Wolcott-Green et al. (2011). We do not invoke an external LW background in addition to in-situ LW radiation from the protostars [^2] This approximation is supported by the build-up of local opacity to external LW radiation due to relic H ii regions in the early intergalactic medium (Johnson et al., 2007)..
@@ 3 Results
Figure 1 plots the density-weighted projections of the number density of the gas at the end of the simulations. The projection window is $0.01\,\rm{pc}$ wide. It is centered on the isolated star in the HD and RHD runs, and uses the center of mass of all stars in the MHD and RMHD runs. We plot the density-weighted projections of the gas temperature for the four simulations in Figure 2, and phase diagrams for the entire cloud in Figure 3. We see from Figure 1 and Figure 2 that the central star in the HD and RHD runs is fed by a well-defined, thermal pressure-supported accretion disc that remains stable against fragmentation for the entirety of the simulation. The high temperatures in these runs is a result of dissociation of H2 by shocks, a process that is only resolved in simulations with sufficiently high Jeans resolution (e.g., Turk et al., 2012; Sharda et al., 2021; Sharda & Menon, 2024). The key difference between the HD and RHD runs is the presence of cooler, H2-dominated gas near the star in the latter, where H2 dissociation is prevented due to radiation pressure slowing down accretion shocks (see also, Sharda & Menon 2024). Given the relatively high accretion rates, the star is unable to contract and produce significant ionizing (EUV) photons that can ionize H or H2 (e.g., Omukai & Palla, 2001; Hosokawa & Omukai, 2009; Haemmerlé et al., 2018). In contrast to HD and RHD simulations, the gas is cooler in MHD and RMHD runs across the entire central envelope. Magnetic fields suppress gravitational collapse, thereby also reducing the rate of compressional heating. Reduced heating allows the gas to remain molecular, and H2 cooling further ensures gas temperatures remain low. These figures show, for the first time, the simultaneous effects of magnetic fields, and ionizing as well as dissociating radiation feedback during the evolution of primordial clouds.
@@
We see from Figure 4 that the amount of mass reaching the envelope initially increases in the MHD and RMHD runs, but then turns over and starts to decline. In contrast, the rate of mass transfer from the envelope to the accretion disk is initially quite fast in the HD and RHD runs, resulting in a decline of mass in the envelope and buildup of mass in the disk. This mass is consequently accreted by the star at a high rate. Since the mass of the star is initially small, the ratio $M_{\rm{disk}}/M_{\star}$ is high (see also Sharda & Menon, 2024, fig. 8). We learn that magnetic fields limit the amount of gas infalling onto the envelope at later stages by acting against gravity, leading to mass depletion within the accretion disk. Fragmentation events correlate with a large buildup of mass in the accretion disc such that the ratio of mass in the disc to the protostellar mass increases far beyond unity. The presence of multiple stars reduces the mass available within the disk of the primary star in the MHD and RMHD runs, whereas the disk mass in the HD and RHD runs continues to build up.
Figure 4: Top panel: evolution of mass enclosed within an envelope of radius $0.01\,\rm{pc}$, centered at the location of the most massive star in each simulation. The MHD and RMHD runs fragment $1500\,\rm{yr}$ and $2235\,\rm{yr}$ after the formation of the first star, which is demarcated by the onset of dashed orange and purple curves, respectively. Bottom panel: mass enclosed within a disk of radius $500\,\rm{au}$ and height $50\,\rm{au}$ (from the midplane) around the most massive star. The mass reservoir that can be accreted onto the central star in the MHD and RMHD runs eventually decreases as magnetic fields suppress gravitational collapse.
@@
The top panel of Figure 5 shows that the accretion rates (of the most massive star) are of the order of $0.03\,\rm{M_{\odot}}\,yr^{-1}$ in the first $1000\,\rm{yr}$, and decline as time progresses. Interestingly, the accretion rates in the MHD and RMHD runs remain systematically lower than the HD and RHD runs for the most part. The most massive star in the MHD run even stops accreting for a considerable period of time. The key impact of lower accretion rates in the MHD and RMHD runs is that the most massive stars only build up half as much mass as the HD and RHD runs within the same time period. We show this in the bottom panel of Figure 5. The mass of the (isolated) star at the end of the simulations in the HD and RHD runs is $127\,\rm{M_{\odot}}$ and $120\,\rm{M_{\odot}}$, respectively. On the other hand, the mass of the most massive star in the MHD and RMHD runs is $48\,\rm{M_{\odot}}$ and $67\,\rm{M_{\odot}}$, respectively. However, the integrated star formation efficiency (SFE; defined as the cumulative mass of all stars normalized by the initial cloud mass; solid orange and purple curves in the bottom panel of Figure 5) is lower by 30 per cent in the RMHD run as compared to all the control runs.
There is a subtle but important difference between magnetic fields suppressing mass transport, and fragmentation-induced starvation. The former affects both the mass of the primary star and the integrated SFE (dashed and solid lines in the bottom panel of Figure 5), whereas the latter affects the mass of the primary star mass without necessarily affecting the integrated SFE. Thus, the ‘mass-limiting’ effects of the runs including magnetic fields are not simply due to fragmentation-induced starvation. The slower mass growth and consequently lower star formation efficiency in the presence of magnetic fields has also been observed in simulations of Population I massive star formation, although the differences arise much later on, and are smaller in magnitude (e.g., Rosen & Krumholz, 2020; Kim et al., 2021). Our findings are also in qualitative agreement with Sharda & Menon (2024), where the authors simulated Pop III star formation with magnetic fields and only ionizing feedback.
@@ 4 Evolution on the main sequence
We plot the resulting mass growth from the three cases as blue, orange and green curves in Figure 6, respectively. Despite different accretion histories, the star in all the three cases reaches ZAMS around $13000\,\rm{yr}$ (marked by diamonds in Figure 6) with stellar mass between $90-100\,\rm{M_{\odot}}$. However, the main sequence evolution is significantly different, and the TAMS mass spans a large range. The stars are expected to turn into black holes soon after TAMS (denoted by circles in Figure 6). The star in Case A oscillates between two branches, leading to wiggles in the HR diagram, because its evolution is sensitive to atmospheric parameters (Herrington et al., 2023). This phenomenon only occurs for Case A because the accretion rate is close to the critical value of $0.01\,\rm{M_{\odot}}$ which separates the blue and red branches of supermassive stellar evolution (Hosokawa & Omukai, 2009; Haemmerlé et al., 2018). The period we simulate with different physics, although short, matters for the final fate of the star. To show this, we use Case C to extrapolate the growth of the star in the RHD simulation (dashed green curve in the left panel of Figure 6). Due to differences in accretion histories in the first $5000\,\rm{yr}$, this star ends into the regime of PISNe, rather than ending its life as a black hole. This is unlikely to occur in reality since magnetic fields are excluded in the RHD simulation, but it shows that omitting key physical processes even during the early stages of evolution can profoundly impact the ultimate fate of Pop III stars.
Given the high stellar effective temperatures and luminosities on the main sequence (right panel of Figure 6), radiation feedback combined with competitive accretion due to fragmentation will potentially limit the final mass to $80\,\rm{M_{\odot}}\leq M_{\star}\lesssim 600\,\rm{M_{\odot}}$, making Case A highly unlikely. The magnetic field strength is expected to remain significant beyond the period we simulate here due to a mixture of turbulent and mean field dynamos (Liao et al., 2021; Sharda et al., 2021), and it can provide additional negative feedback by suppressing accretion onto the star, or generating outflows (Machida & Doi, 2013). This will enhance feedback effects and further limit the maximum stellar mass, such that the actual stellar mass is closer to $80\,\rm{M_{\odot}}$ than $600\,\rm{M_{\odot}}$.
@@ Acknowledgements
We thank Zoltan Haiman, Mark Krumholz, Mariska Kriek and Matthieu Schaller for useful discussions. We also thank SURF for the support in using the Dutch National Supercomputer Snellius. PS is supported by the Leiden University Oort Fellowship and the International Astronomical Union – Gruber Foundation (TGF) Fellowship. SHM and BB acknowledge support through NASA grant No. 80NSSC20K0500 and National Science Foundation (NSF) grant AST-2009679, and the CCA at the Flatiron Institute. RG is supported by the Society of Science Fellowship provided by the University of Notre Dame. BB is grateful for generous support by the David and Lucile Packard Foundation and Alfred P. Sloan Foundation. BDW acknowledges support from NSF grant AAG-2106575. The simulations and data analyses presented in this work used high-performance computing resources provided by the Dutch National Supercomputing Facility SURF via project grant EINF-8292 on Snellius, as well as high-performance computing resources provided by the Centre for Computational Astrophysics (CCA) at the Flatiron Institute, and the Australian National Computational Infrastructure (NCI) through project jh2 in the framework of the National Computational Merit Allocation Scheme and the Australian National University (ANU) Allocation Scheme. This work was performed in part at Aspen Center for Physics, which is supported by National Science Foundation grant PHY-2210452. We acknowledge using the following softwares: FLASH (Fryxell et al., 2000; Dubey et al., 2013), VETTAM (Menon et al., 2022), KROME (Grassi et al., 2014), petsc (Balay et al., 1997; Zhang et al., 2022), Astropy (Astropy Collaboration et al., 2013, 2018, 2022), Numpy (Oliphant, 2006; Harris et al., 2020), Scipy (Virtanen et al., 2020), Matplotlib (Hunter, 2007), yt (Turk et al., 2011), cmasher (Van der Velden, 2020). This research has made extensive use of NASA’s Astrophysics Data System Bibliographic Services (ADS).

### v2

@@ 1 Introduction
However, radiation-magnetohydrodynamics simulations in the era of Pop III star formation remain largely absent. Both magnetic fields and protostellar radiation feedback are critical ingredients that influence (massive) star formation across all metallicities (e.g., Tanaka et al., 2018; Chon et al., 2021, 2024; Sharda & Krumholz, 2022). Therefore, conclusions from prior numerical work aimed at deriving the masses of Pop III stars are likely subject to major uncertainties since they either exclude magnetic fields (e.g., Hosokawa et al., 2011, 2016; Sugimura et al., 2020; Jaura et al., 2022) or radiation feedback (e.g., Turk et al., 2012; Sharda et al., 2020; Prole et al., 2022; Saad et al., 2022; Sadanari et al., 2023). In this work, we use the POPSICLE project (Sharda et al. in prep.) to extend the recent RMHD simulations of Pop III star formation by Sharda & Menon (2024) to include far-UV (FUV) molecule-dissociating radiation feedback in addition to extreme-UV (EUV) ionizing feedback. Our aim in this Letter is to show that magnetic fields significantly limit the mass growth of massive Pop III stars, even before radiative feedback becomes dominant. We arrange the remainder of the paper as follows: Section 2 summarizes the setup we use to develop and run the simulations, and Section 3 describes the results. In Section 4, we look at the main sequence evolution of simulated stars using Modules for Experiments in Stellar Astrophysics (MESA, Paxton et al., 2011, 2013, 2015, 2018, 2019). Finally, we summarize in Section 5.
@@ 2 POPSICLE Simulations
We briefly describe the POPSICLE [^1] POPulation II/III Simulations Including Chemistry, Luminosity, and Electromagnetism. project setup employed in this work, and point the reader to Sharda & Menon (2024) for details of the numerical implementation. We use a custom version of the adaptive mesh refinement (AMR) code FLASH (Fryxell et al., 2000; Dubey et al., 2008) which employs an approximate Riemann solver for magnetohydrodynamics (Waagan, 2009; Waagan et al., 2011). We include non-equilibrium primordial chemistry (with deuterium) from the KROME astrochemistry package (Grassi et al., 2014). We use the VETTAM radiation hydrodynamics scheme, which uses a non-local variable Eddington tensor (VET) closure obtained with a ray-trace solve to close the radiation moment equations (Menon et al., 2022). We use the GENEVA Pop III protostellar model grid to evolve radiative properties of the protostars as a function of their mass and accretion rates (Haemmerlé et al., 2016, 2018). The protostars are represented by sink particles in the simulation, following the criteria described in Federrath et al. (2010b, 2011).
Population III protostars contract and produce significant radiation once accretion rates drop below $0.01\,\rm{M_{\odot}\,yr^{-1}}$ (Omukai & Palla, 2001, 2003). We consider the ionization of H and H2 due to extreme-UV (EUV) photons released from the protostar(s) with energies upwards of $13.6\,\rm{eV}$ (upwards of $15.2\,\rm{eV}$ for H2). The key physics we add to the simulations by Sharda & Menon (2024) is the dissociation of H2 by FUV photons in the Lyman-Werner (LW) band, between $11.2-13.6\,\rm{eV}$, which can significantly affect the thermodynamic state of dense gas in the vicinity of the stars. Importantly, we consider both self-shielding of H2 as well as cross shielding by H (Wolcott-Green & Haiman, 2011); the latter has been ignored in previous works. We adopt the fitting functions for self- and cross-shielding of H2 from Wolcott-Green et al. (2011). We do not invoke an external LW background in addition to in-situ LW radiation from the protostars [^2] This approximation is supported by the build-up of local opacity to external LW radiation due to relic H ii regions in the early intergalactic medium (Johnson et al., 2007)..
@@ 3 Results
Figure 1 plots the density-weighted projections of the number density of the gas at the end of the simulations. The projection window is $0.01\,\rm{pc}$ wide. It is centered on the isolated star in the HD and RHD runs, and uses the center of mass of all stars in the MHD and RMHD runs. We plot the density-weighted projections of the gas temperature for the four simulations in Figure 2, and phase diagrams for the entire cloud in Figure 3. We see from Figure 1 and Figure 2 that the central star in the HD and RHD runs is fed by a well-defined, thermal pressure-supported accretion disc that remains stable against fragmentation for the entirety of the simulation. The high temperatures in these runs is a result of dissociation of H2 by shocks, a process that is only resolved in simulations with sufficiently high Jeans resolution (e.g., Turk et al., 2012; Sharda et al., 2021; Sharda & Menon, 2024). The key difference between the HD and RHD runs is the presence of cooler, H2-dominated gas near the star in the latter, where H2 dissociation is prevented due to radiation pressure slowing down accretion shocks (see also, Sharda & Menon 2024). Given the relatively high accretion rates, the star is unable to contract and produce significant ionizing (EUV) photons that can ionize H or H2 (e.g., Omukai & Palla, 2001; Hosokawa & Omukai, 2009; Haemmerlé et al., 2018). In contrast to HD and RHD simulations, the gas is cooler in MHD and RMHD runs across the entire central envelope. Magnetic fields suppress gravitational collapse, thereby also reducing the rate of compressional heating. Reduced heating allows the gas to remain molecular, and H2 cooling further ensures gas temperatures remain low.
These figures show, for the first time, the simultaneous effects of magnetic fields, and ionizing as well as dissociating radiation feedback during the evolution of primordial clouds. Note that we do not include the effects of radiative heating caused by accretion luminosity, which could lead to higher gas temperatures close to the protostars, thereby reducing the accretion rates and limiting protostellar mass growth (e.g., Smith et al., 2011; Wollenberg et al., 2020). However, it cannot hinder accretion for long time periods since low accretion rates in turn lead to lower accretion heating. Further, accretion luminosity heating only becomes significant when the opacity of (primordial) gas is large, which occurs for $T>5000\,\rm{K}$ at the densities we resolve (Mayer & Duschl, 2005, Table E3). Given that magnetic fields lead to lower average gas temperatures, we expect accretion luminosity to be less important in the MHD and RMHD simulations as compared to HD and RHD simulations.
@@
We see from Figure 4 that the amount of mass reaching the envelope initially increases in the MHD and RMHD runs, but then turns over and starts to decline. Despite the envelope and the disc containing larger masses early on in the MHD and RMHD runs, the corresponding accretion rates onto the protostars are lower (see Figure 5). Close to the protostar, the magnetic field is sub-Alfvénic (plasma $\beta<1$), significantly inhibiting mass transport. In contrast, the rate of mass transfer from the envelope to the accretion disc is initially quite fast in the HD and RHD runs, resulting in a decline of mass in the envelope and buildup of mass in the disc. This mass is consequently accreted by the star at a high rate. Fragmentation events correlate with a large buildup of mass in the accretion disc such that the ratio of mass in the disc to the protostellar mass increases far beyond unity. The presence of multiple stars reduces the mass available within the disc of the primary star in the MHD and RMHD runs, whereas the disc mass in the HD and RHD runs continues to build up.
Figure 4: Top panel: evolution of mass enclosed within an envelope of radius $0.01\,\rm{pc}$, centered at the location of the most massive star in each simulation. The MHD and RMHD runs fragment $1500\,\rm{yr}$ and $2235\,\rm{yr}$ after the formation of the first star, which is demarcated by the onset of dashed orange and purple curves, respectively. Bottom panel: mass enclosed within a disc of radius $500\,\rm{au}$ and height $50\,\rm{au}$ (from the midplane) around the most massive star. The mass reservoir that can be accreted onto the central star in the MHD and RMHD runs eventually decreases as magnetic fields suppress gravitational collapse.
@@
The top panel of Figure 5 shows that the accretion rates (of the most massive star) are of the order of $0.03\,\rm{M_{\odot}}\,yr^{-1}$ in the first $1000\,\rm{yr}$, and decline as time progresses. The accretion rates in the MHD and RMHD runs remain systematically lower than the HD and RHD runs for the most part, initially due to strong magnetic fields inhibiting accretion, and later due to fragmentation. The key impact of lower accretion rates in the MHD and RMHD runs is that the most massive stars only build up half as much mass as the HD and RHD runs within the same time period. We show this in the bottom panel of Figure 5. The mass of the (isolated) star at the end of the simulations in the HD and RHD runs is $127\,\rm{M_{\odot}}$ and $120\,\rm{M_{\odot}}$, respectively. On the other hand, the mass of the most massive star in the MHD and RMHD runs is $48\,\rm{M_{\odot}}$ and $67\,\rm{M_{\odot}}$, respectively. However, the integrated star formation efficiency (SFE; defined as the cumulative mass of all stars normalized by the initial cloud mass; solid orange and purple curves in the bottom panel of Figure 5) is lower by 30 per cent in the RMHD run as compared to all the control runs.
There is a subtle but important difference between magnetic fields suppressing mass transport, and fragmentation-induced starvation. The former turns on earlier, and affects both the mass of the primary star and the integrated SFE (dashed and solid lines in the bottom panel of Figure 5), whereas the latter emerges later, and affects the mass of the primary star mass without necessarily affecting the integrated SFE. Thus, the ‘mass-limiting’ effects of the runs including magnetic fields are not simply due to fragmentation-induced starvation. The slower mass growth and consequently lower star formation efficiency in the presence of magnetic fields has also been observed in simulations of Population I massive star formation, although the differences arise much later on, and are smaller in magnitude (e.g., Rosen & Krumholz, 2020; Kim et al., 2021). Our findings are also in qualitative agreement with Sharda & Menon (2024), where the authors simulated Pop III star formation with magnetic fields and only ionizing feedback.
@@ 4 Evolution on the main sequence
We plot the resulting mass growth from the three cases as blue, orange and green curves in Figure 6, respectively. Despite different accretion histories, the star in all the three cases reaches ZAMS around $13000\,\rm{yr}$ (marked by diamonds in Figure 6) with stellar mass between $90-100\,\rm{M_{\odot}}$. However, the main sequence evolution is significantly different, and the TAMS mass spans a large range. The stars are expected to turn into black holes soon after TAMS (denoted by circles in Figure 6). The star in Case A oscillates between two branches, leading to wiggles in the HR diagram, because its evolution is sensitive to atmospheric parameters (Herrington et al., 2023). This phenomenon only occurs for Case A because the accretion rate is close to the critical value of $0.01\,\rm{M_{\odot}}\,yr^{-1}$ which separates the blue and red branches of supermassive stellar evolution (Hosokawa & Omukai, 2009; Haemmerlé et al., 2018). The period we simulate with different physics, although short, matters for the final fate of the star. To show this, we use Case C to extrapolate the growth of the star in the RHD simulation (dashed green curve in the left panel of Figure 6). Due to differences in accretion histories in the first $5000\,\rm{yr}$, this star ends into the regime of PISNe, rather than ending its life as a black hole. This is unlikely to occur in reality since magnetic fields are excluded in the RHD simulation, but it shows that omitting key physical processes even during the early stages of evolution can profoundly impact the ultimate fate of Pop III stars.
Given the high stellar effective temperatures and luminosities on the main sequence (right panel of Figure 6), radiation feedback combined with competitive accretion due to fragmentation will potentially limit the final mass to $80\,\rm{M_{\odot}}\leq M_{\star}\lesssim 600\,\rm{M_{\odot}}$, making Case A highly unlikely. The magnetic field strength is expected to remain significant beyond the period we simulate here due to a mixture of turbulent and mean field dynamos (Liao et al., 2021; Sharda et al., 2021), and it can provide additional negative feedback by suppressing accretion onto the star, or generating outflows (Machida & Doi, 2013). Together with the effects of heating due to accretion luminosity we discuss in Section 3, this will enhance feedback effects and further limit the maximum stellar mass, such that the actual stellar mass is closer to $80\,\rm{M_{\odot}}$ than $600\,\rm{M_{\odot}}$.
@@ Acknowledgements
We thank Zoltan Haiman, Simon Glover, Mark Krumholz, Mariska Kriek and Matthieu Schaller for useful discussions, and the anonymous referee for their feedback that helped improve this work. We also thank SURF for the support in using the Dutch National Supercomputer Snellius. PS is supported by the Leiden University Oort Fellowship and the International Astronomical Union – Gruber Foundation (TGF) Fellowship. SHM and BB acknowledge support through NASA grant No. 80NSSC20K0500 and National Science Foundation (NSF) grant AST-2009679, and the CCA at the Flatiron Institute. RG is supported by the Society of Science Fellowship provided by the University of Notre Dame. BB is grateful for generous support by the David and Lucile Packard Foundation and Alfred P. Sloan Foundation. BDW acknowledges support from NSF grant AAG-2106575. The simulations and data analyses presented in this work used high-performance computing resources provided by the Dutch National Supercomputing Facility SURF via project grant EINF-8292 on Snellius, as well as high-performance computing resources provided by the Centre for Computational Astrophysics (CCA) at the Flatiron Institute, and the Australian National Computational Infrastructure (NCI) through project jh2 in the framework of the National Computational Merit Allocation Scheme and the Australian National University (ANU) Allocation Scheme. This work was performed in part at Aspen Center for Physics, which is supported by National Science Foundation grant PHY-2210452. We acknowledge using the following softwares: FLASH (Fryxell et al., 2000; Dubey et al., 2013), VETTAM (Menon et al., 2022), KROME (Grassi et al., 2014), petsc (Balay et al., 1997; Zhang et al., 2022), Astropy (Astropy Collaboration et al., 2013, 2018, 2022), Numpy (Oliphant, 2006; Harris et al., 2020), Scipy (Virtanen et al., 2020), Matplotlib (Hunter, 2007), yt (Turk et al., 2011), cmasher (Van der Velden, 2020). This research has made extensive use of NASA’s Astrophysics Data System Bibliographic Services (ADS).